You are given the production function: Q(K,L)=10KαLβ

What is the marginal rate of technical substitution (MRTS)? Simplify fully


1
Expert's answer
2021-05-17T10:39:57-0400

Solution:

Marginal Rate of Technical Substitution (MRTS) is the rate at which a firm can substitute capital with labor. MRTS equals the slope of an isoquant, which is a curve that represents the different combinations of multiple factors of production, that is labor and capital that yields the same total production. MRTS tells us the rate at which a producer is willing to switch one input (factor of production) with another.

MRTS is calculated as follows:


MRTS = "\frac{Change\; in \;Capital}{Change\; in\; Labor} = \frac{MP_{K} }{MP_{L} }"


MRTS(K,L) ="\frac{MP_{K} }{MP_{L} }"


MPL = "\\frac{\u03b4Q}{\u03b4L} = 10\u03b2K^{\u03b1}L ^{\u03b2\u22121}\n\u200b"


MPK = "\\frac{\u03b4Q}{\u03b4K} = 10\u03b1K^{\u03b1\u22121}L ^\u03b2\n \n\u200b"


MRTS = "\frac{10βK^{α}L ^{β−1} }{10αK^{α−1}L ^β } = \frac{βK}{αL }"


MRTS ="\frac{βK}{αL }"